# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

### v1

@@ Abstract
Stably stratified shear flows, where the base velocity shear is quasi-continuously forced externally, can arise in many geophysically and environmentally relevant circumstances, for example at estuarine outflows. It is important to determine the emergent dynamics of the ensuing statistically steady stratified turbulence. We investigate this phenomenon in a series of three-dimensional direct numerical simulations using spectral element methods. We force the flow to relax back towards vertical hyperbolic tangent profiles of streamwise velocity and buoyancy, with characteristic half-depth $d_{0}$, half-velocity jump $U_{0}$, and half-buoyancy jump $B_{0}$, with a relaxation time $t_{r}=100\tau_{\textrm{adv}}$ where $\tau_{\textrm{adv}}:=d_{0}/U_{0}$. We consider computational domains of vertical extent $L_{z}=48$ (scaled with $d_{0}$) with a range of (scaled) horizontal extents $16\leq L_{\mathrm{h}}\leq 512$. We simulate a fluid with Prandtl number $\mathrm{Pr}:=\nu/\kappa=1$, where $\nu$ is the kinematic viscosity and $\kappa$ is the buoyancy diffusivity, and set the initial bulk Reynolds number $\mathrm{Re}_{0}:=U_{0}d_{0}/\nu=50$ and initial bulk Richardson number $\mathrm{Ri}_{0}:=B_{0}d_{0}/U_{0}^{2}=1/80$ (corresponding to the initial minimum gradient Richardson number). At these parameters, the flow is initially (vigorously) unstable to a primary Kelvin-Helmholtz instability. We simulate the continuously forced flow over about $5000\tau_{\textrm{adv}}$, and investigate the dynamically emergent length scales and turbulence properties of the statistically stationary flow, in particular the local turbulent flux coefficient $\Gamma_{\chi}$, as well as streamwise and spanwise Fourier spectra. We find that the shear layer half depth converges to $d\approx 8$ (i.e., $\Lambda_{z}\approx 16$) with markedly increased minimum gradient Richardson number ($\sim 0.1$) and $\mathrm{Re}\approx 400$ with associated convergent vertical mixing properties only for $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. However, emergent dominant yet large-scale spanwise or streamwise flow structures appear to extend up to (scaled) $\Lambda_{y}\approx 50$ or even $\Lambda_{x}\approx 115$, respectively, and thus may only be fully resolved for horizontal domains of extent $L_{\mathrm{h}}\gtrsim 256$. Our observations demonstrate the marked anisotropy of characteristic emergent length scales, even for such ‘weakly stratified’ forced shear flows, and are consistent with the possibility that an ‘imprint’ of the primary linear instability even continues to survive in such vigorously turbulent flows, as the streamwise flow structures have a scale consistent with the most unstable KHI for the converged deepened turbulent shear layer.
@@ 1 Introduction
Stably stratified (vertical) shear flows, where both the background fluid buoyancy (i.e. the appropriatedly scaled negative density perturbation) and horizontal flow velocity vary with height, are ubiquitous in geophysical contexts. There has been a very large body of work considering the ways in which such flows behave, (as evidenced by a sequence of reviews such as Fernando (1991); Peltier & Caulfield (2003); Ivey et al. (2008); Caulfield (2020) and Caulfield (2021)) with significant focus on the (possible) existence of turbulence extracting and converting the kinetic energy in the background shear and the associated enhanced (irreversible) turbulent mixing. Understanding (and parameterising) such turbulent mixing of stratified fluids is a key challenge in geophysical flows, as the transport of momentum, heat and other scalars (such as dissolved gases, pollution and microplastics for example) is both a crucial process and a phenomenon of great uncertainty for the description of the world’s climate and environment (see for example the reviews of Wunsch & Ferrari (2004); Ferrari & Wunsch (2009); Gregg et al. (2018)).
@@
Figure 1: Fundamental configuration of forced stratified shear flows, with an initial buoyancy profile $b_{0}$ that is statically stable, and an imposed initial velocity profile $u_{x,0}$ which may induce flow instability leading to a transtion to turbulence. Right: A continuous forcing $f_{\Phi}$ injects the required energy to sustain this turbulence, ensuring statistically stationary dynamics.
@@
In essence, as shown schematically in figure 1, stably stratified shear flows are characterised by a competition between a stabilising buoyancy and a de-stabilising velocity (or shear) profile. However, understanding fundamental aspects of this deceptively simple set-up is exceptionally complicated. Considering initial value problems of (initially) laminar velocity and density profiles, it is well-known that such flows can be prone to a range of primary flow instabilities (see the review of Caulfield (2021)) that effectively rearrange the strip of spanwise vorticity into trains of elliptical vortices (such as the classic Kelvin-Helmholtz ‘billows’, i.e. the saturated manifestation of the Kelvin-Helmholtz Instability (KHI)) which themselves are prone (at least for sufficiently high Reynolds number) to a ‘zoo’ of secondary instabilities (a nomenclature proposed by Mashayek & Peltier (2012a, b)) which trigger turbulence transition and hence significantly enhanced mixing and dissipation. As a (perhaps tenuous) paradigm for mixing induced by breaking internal waves in the atmosphere and ocean, direct numerical simulations of such transient mixing events have been widely conducted. There is then an inevitable set of trade-offs in how best to utilise finite computational resources. The need to simulate the flow at as high a Reynolds number as possible (defined in terms of the shear layer half depth $d_{0}$ and half velocity difference $U_{0}$ as shown in figure 1) to ensure sufficiently realistic turbulence must be balanced against the need for the computational domain to be sufficiently large to allow realistic flow structures, as well as the need to compute for sufficiently long times to capture appropriately the full life cycle of such mixing events.
@@
Although such mixing events can extend for a significant length of time (particularly when the flow is prone primarily to the so-called ‘Holmboe Wave Instability’ (HWI) (Holmboe, 1962), as discussed in detail by Salehipour et al. (2016, 2018)), there will always be an inevitable decay, unless the underlying driving mechanism (the shear) is replenished. The particular life cycle of such mixing events can be strongly sensitive to initial conditions, as demonstrated conclusively by Liu et al. (2022), due to competition between different members of the secondary instability ‘zoo’, especially involving the subharmonic ‘pairing’ instability of neighbouring primary Kelvin-Helmholtz billows. As demonstrated by Mashayek & Peltier (2013), sufficiently strong stratification can disrupt and hence suppress such pairing (or more precisely ‘merging’) events due to the energetic costs of the inherent vertical motions. Such suppression is often cited as a reason to restrict computational domains considering such mixing events to one (or at most two) wavelengths of the primary instabilities, allowing the simulations to be conducted at sufficiently high Reynolds number for the ensuing mixing not to be (excessively) dominated by viscous and diffusive effects.
@@
However, restricting the horizontal extent of the computational domain inevitably affects the flow dynamics also. As shown by Scinocca & Ford (2000), rich dynamics can occur in longer streamwise domains, where primary instabilities with close wavelengths (and linear growth rates) can compete as they grow to finite amplitude if the (conventional) imposition of periodic streamwise boundary conditions does not quantise the possible instabilities too severely. Analogous issues arise with the spanwise extent of the computational extent. Many studies consider relatively narrow computational domains, in the sense that the spanwise extent is (often significantly) smaller than the characteristic (streamwise) wavelength of the primary instability. Such domains allow many of the (essentially local) secondary instabilities to develop and hence trigger turbulence transition. However, as can be straightforwardly observed in sufficiently wide tilting tank experiments (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and sufficiently wide computational domains, as clearly demonstrated by Fritts et al. (2022a, b), inherently three-dimensional ‘knots’, ‘tubes’, and billow ‘defects’ can develop in the spanwise direction on scales of the order (but typically larger) of the primary instability’s (streamwise) wavelength.
Although such initial-value-problem mixing events are undoubtedly of geophysical interest and indeed demonstrate that the chosen size of the computational domain can affect the emergent flow structures, it is exceptionally difficult to determine to what extent the size of the considered flow domain affects the ensuing stratified turbulence (and associated quantities of interest, such as the mixing), due not least to the inherent transient nature of the mixing event. However, it is important to remember that such initial-value-problem mixing events are really just one end member class of the possible flows of geophysical interest.
@@
The other obvious end member class is the class of continuously forced flows, where the ‘background’ velocity and density (or equivalently buoyancy) profiles are driven by some external, quasi-steady forcing. Possible candidate mechanisms for such geophysically relevant forcings include wind, solar radiation and resulting evaporation at the surface of the ocean (Thorpe, 2005), tidal forcing (Laurent et al., 2002), active matter living in water (Castro et al., 2022), or continuous outflows from rivers (Uncles & Mitchell, 2011). While this list is not meant to be complete in any sense, it illustrates that quasi-steady forcings do occur in geophysically relevant situations. An (artificial) volumetric forcing is a particularly convenient way to mimic these natural complex mechanisms in a simplified manner. As visualised on the right side of figure 1, such a forcing may be defined to relax the local profiles.
@@
Such a forcing was used by Smith et al. (2021), who demonstrated that, after an initial transient where primary instabilities (either KHI or HWI) develop and (inevitably) break down, the ensuing turbulence could be sustained over arbitrarily long times, thus enabling statistically steady statistics of the flow to be calculated. There is a clear attraction to considering such forced flows due principally to the inherent removal of the confounding effects of the life cycle of the mixing event from the turbulence statistics. Therefore, such flows seem the natural test-bed to consider the effect of the size of the computational domain on the convergence (or not) of such statistics, in an attempt to identify how large a computational domain needs to be for the imposition of horizontal periodic boundary conditions to become insignificant. However, as in the initial value problem flows discussed above, Smith et al. (2021) also considered relatively small computational domains (in terms of the initial shear layer half-depth $d_{0}$), and so it is (again) not at all clear whether or not the flow remained unaffected by the computational domain size. An equivalent question to ask is what are the (unconstrained) emergent length scales of a forced stratified shear flow, and answering that (open) question is the central objective of this paper.
@@
To address this question, the rest of the paper is organised as follows. In section 2 we present our numerical approach before we list the range of considered computational domains in section 3. For simplicity, we focus on a relatively weakly stratified flow, prone to primary KHI, that initially ‘rolls up’ into a train of Kelvin-Helmholtz billows. We identify some interesting early-time dynamics that may possibly be affected by diffusion. However, it is not this state that is of principal interest, but rather the long-time statistically steady (yet sheared) turbulent state. We present the results of those simulations, identifying the (strongly anisotropic) length scales which emerge in sufficiently large flow domains. We also demonstrate conclusively that key statistics of the flow, including in particular those related to irreversible mixing, are sensitive to the size of the computational domain. We remark that for convergent statistics, perhaps surprisingly, the flow domain needs to be extraordinarily ‘large’ or extended compared to the initial shear-layer half-depth (i.e. of order $100$ times larger). Finally, in section 4 we draw our conclusions, in particular discussing the implications of our results for future research.
@@ 2.1 Governing equations
where $t_{\textrm{r}}$ is the response time while $u_{x,0}$ and $b_{0}$ represent the initial streamwise velocity and buoyancy base profiles to which the flow relaxes back. In this context, $R_{0}:=d_{0}/\delta_{0}$ defines the ratio of initial interface half-depths (i.e. $\delta_{0}$ represents the dimensional initial buoyancy interface half-depth) with $R_{0}=\sqrt{\mathrm{Pr}}$ following the diffusive arguments presented by Smyth et al. (1988). In essence, these forcing terms are idealized models of geophysically relevant processes that tend to restore the initial profiles $u_{x,0}$ and $b_{0}$, and thus sustain turbulence over arbitrarily long times.
Hence, the governing equations (2) – (4) are fully specified via four control parameters: $\mathrm{Pr}$, $\mathrm{Re}_{0}$, $\mathrm{Ri}_{0}$, and $t_{\textrm{r}}$. The associated energy equations are provided in appendix A.
@@ 2.2 Domain, boundary and initial conditions, and numerical code
As indicated by figure 1, the streamwise, spanwise, and vertical extents of the Cartesian numerical domain are $L_{x}$, $L_{y}$, and $L_{z}$, respectively. Both the midpoint of the shear layer and the midpoint of the buoyancy distribution are located at midplane, $z=0$, with a horizontal cross-section $A:=L_{x}\times L_{y}$. We consider a horizontally periodic domain where any quantity
@@
The random fluctuations $-10^{-3}\leq\Upsilon\left(\bm{x}\right)\leq 10^{-3}$ ‘seed’ the development of primary instabilities.
We solve the coupled governing equations (2) – (4), subject to these boundary and initial conditions, using the GPU-accelerated spectral element solver NekRS (Fischer, 1997; Scheel et al., 2013; Fischer et al., 2022). As shown in more detail in appendix B, spectral element methods can accommodate different required spatial resolutions across the domain and are thus perfectly suited to resolve shear flows efficiently. This is particularly important given degrees of freedom of up to $N_{\textrm{dof}}\approx 3N_{\textrm{e}}N^{3}\sim\mathcal{O}\left(10^{9}\right)$ in our present study.
@@
$16^{2}$ 00 $9^{2}$ 00 $2.66\pm 0.16$ $2.80\pm 0.14$ $0.95\pm 0.04$ $133\pm\hphantom{0}8$ $0.0333\pm 0.0020$
$32^{2}$ 00 $18^{2}$ 00 $5.35\pm 0.34$ $5.54\pm 0.35$ $0.96\pm 0.02$ $267\pm 17$ $0.0668\pm 0.0043$
$64^{2}$ 00 $36^{2}$ 00 $7.55\pm 0.47$ $7.75\pm 0.36$ $0.97\pm 0.03$ $377\pm 24$ $0.0943\pm 0.0059$
$96^{2}$ 00 $54^{2}$ 00 $7.92\pm 0.53$ $8.27\pm 0.56$ $0.96\pm 0.02$ $396\pm 26$ $0.0990\pm 0.0066$
$128^{2}$ 00 $72^{2}$ 00 $8.05\pm 0.42$ $8.41\pm 0.45$ $0.96\pm 0.01$ $403\pm 21$ $0.1007\pm 0.0053$
$256^{2}$ 00 $144^{2}$ 00 $7.90\pm 0.18$ $8.22\pm 0.17$ $0.96\pm 0.01$ $395\pm\hphantom{0}9$ $0.0987\pm 0.0022$
$512^{2}$ 00 $288^{2}$ 00 $7.73\pm 0.08$ $8.06\pm 0.07$ $0.96\pm 0.00$ $387\pm\hphantom{0}4$ $0.0967\pm 0.0010$
@@
Table 1: Simulation parameters. The Prandtl number $\mathrm{Pr}=1$, initial bulk Reynolds number $\mathrm{Re}_{0}=50$, initial bulk Richardson number $\mathrm{Ri}_{0}=0.0125$, response time $t_{\textrm{r}}=100$, and initial ratio of interface (half) thicknesses $R_{0}=1$ in a horizontally periodic domain. In the vertical direction, the domain has an aspect ratio $L_{z}=48$ spanned by $N_{\mathrm{e},z}=18$ non-uniformly distributed spectral elements (see appendix B for more details) together with free-slip and no-flux boundary conditions for the velocity and buoyancy field, respectively. The polynomial order $N=6$. Although the total evolution or run time of each flow $t_{\textrm{evo}}=5,040$, this work focuses on the statistically stationary dynamics during the last $\Delta t=3,000$ only. For each simulation, this table lists the horizontal aspect ratios $L_{x}\times L_{y}$ and corresponding numbers of uniformly distributed spectral elements $N_{\mathrm{e},x}\times N_{\mathrm{e},y}$. Moreover, we include the final and dynamically manifesting shear layer (half) depth $d$ of the streamwise velocity field as well as the buoyancy interface half-depth $\delta$, their ratio $R$, the bulk Reynolds number $\mathrm{Re}$, as well as the bulk Richardson number $\mathrm{Ri}$, listing both temporal means and standard deviations.

### v2

@@ Abstract
Although stably stratified shear flows, where the base velocity shear is quasi-continuously forced externally, arise in many geophysically and environmentally relevant circumstances, the emergent dynamics of their ensuing statistically steady stratified turbulence is still an open question. We address this phenomenon in a series of three-dimensional direct numerical simulations using spectral element methods. We consider a forced, stably stratified shear flow with an initial bulk Reynolds number $\mathrm{Re}_{0}=50$, an initial bulk Richardson number $\mathrm{Ri}_{0}=1/80$ (also corresponding to the initial minimum gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$), and a fluid of Prandtl number $\mathrm{Pr}=1$ in horizontally extended domains. Although the initial configuration is unstable to a primary Kelvin-Helmholtz instability, the ensuing turbulence is sustained by continuously relaxing the resulting flow back towards the initial profiles of streamwise velocity and buoyancy. We study statistical as well as structural aspects of the final statistically steady flows, including the flux coefficient $\Gamma_{\chi}$ and dynamically emergent length scales $\Lambda$ associated with the large-scale dynamics, respectively. Despite the ongoing stirring and mixing, we find that the shear layer half-depth converges to a finite value of $d\approx 8$ (i.e., $\Lambda_{z}\approx 16$) once the horizontal extent of the domain $L_{\mathrm{h}}\gtrsim 96$. While this implies a final $\mathrm{Re}\approx 400$ and $\mathrm{Ri}\approx 0.1$, we hypothesise that such forced flows ‘tune’ themselves eventually to a state of a gradient Richardson number $\mathrm{Ri}_{\textrm{g}}\lesssim 0.2$, consistently with several previous studies. Moreover, provided sufficiently extended domains, we observe the emergence of large-scale flow structures with spanwise $\Lambda_{y}\approx 50$ and streamwise $\Lambda_{x}\lesssim 115$. Clearly, these observations demonstrate the marked anisotropy of characteristic emergent length scales, even for such ‘weakly stratified’ forced shear flows. We conjecture that the actual emergent streamwise structures are a vestigial ‘imprint’ in the sheared turbulent flow of the primary linear instability of the converged deepened turbulent shear layer.
@@ 1 Introduction
Stably stratified (vertical) shear flows, where both the background fluid buoyancy (i.e. the appropriately scaled negative density perturbation) and horizontal flow velocity vary with height, are ubiquitous in geophysical contexts. There has been a very large body of work considering the ways in which such flows behave, (as evidenced by a sequence of reviews such as Fernando (1991); Peltier & Caulfield (2003); Ivey et al. (2008); Caulfield (2020) and Caulfield (2021)) with significant focus on the (possible) existence of turbulence extracting and converting the kinetic energy in the background shear and the associated enhanced (irreversible) turbulent mixing. Understanding (and parameterising) such turbulent mixing of stratified fluids is a key challenge in geophysical flows, as the transport of momentum, heat and other scalars (such as dissolved gases, pollution and microplastics for example) is both a crucial process and a phenomenon of great uncertainty for the description of the world’s climate and environment (see for example the reviews of Wunsch & Ferrari (2004); Ferrari & Wunsch (2009); Gregg et al. (2018)).
@@
Figure 1: Configuration of the forced stratified shear flows considered here. While the initial buoyancy profile $b_{0}$ is statically stable, the imposed initial velocity profile $u_{x,0}$ may induce flow instabilities that lead to a transition to turbulence. A continuous forcing $f_{\Phi}$ may inject the required energy to sustain this turbulence, ensuring statistically stationary dynamics.
@@
In essence, as shown schematically in figure 1, stably stratified shear flows are characterised by a competition between a stabilising buoyancy and a de-stabilising velocity (or shear) profile. However, gaining an understanding of fundamental aspects of this deceptively simple set-up continues to prove exceptionally complicated. Considering initial value problems of (initially) laminar velocity and density profiles, it is well-known that such flows can be prone to a range of primary flow instabilities (see the review of Caulfield (2021)) that effectively rearrange the strip of spanwise vorticity into trains of elliptical vortices (such as the classic Kelvin-Helmholtz ‘billows’, i.e. the saturated manifestation of the Kelvin-Helmholtz Instability (KHI)). As these vortices themselves are prone (at least for sufficiently high Reynolds number) to a ‘zoo’ of secondary instabilities (a nomenclature proposed by Mashayek & Peltier (2012a, b)), turbulence transition is triggered and mixing as well as dissipation are eventually significantly enhanced.
@@
Such mixing events can extend for a significant length of time – particularly when the flow is prone primarily to the so-called ‘Holmboe Wave Instability’ (HWI) (Holmboe, 1962), as discussed in detail by Salehipour et al. (2016, 2018)and can also be accompanied by the formation of larger-scale emergent structures, especially in sufficiently spatially extended flow domains (Watanabe & Nagata, 2021). The particular life cycle of such mixing events can be strongly sensitive to initial conditions, as demonstrated conclusively by Liu et al. (2022), due to competition between different members of the secondary instability ‘zoo’, especially involving the subharmonic ‘pairing’ instability of neighbouring primary Kelvin-Helmholtz billows. As demonstrated by Mashayek & Peltier (2013), sufficiently strong stratification can disrupt and hence suppress such pairing (or more precisely ‘merging’) events due to the energetic costs of the inherent vertical motions. Given only finite computational resources, such suppression is often cited as a reason to restrict (the inevitably finite) domains to one (or at most two) wavelengths of the primary instabilities, allowing to reach higher Reynolds numbers.
@@
However, the horizontal extent of the (numerical or experimental) domain may clearly affect the flow dynamics. As shown by Scinocca & Ford (2000), rich dynamics can occur in longer streamwise domains, where primary instabilities with close wavelengths (and hence close linear growth rates) can compete as they grow to finite amplitude if the (conventional) imposition of periodic streamwise boundary conditions does not quantise the accessible wavelengths of possible instabilities too severely. Analogous issues also arise in the spanwise direction. Many studies consider relatively narrow domains, in the sense that the spanwise extent is (often significantly) smaller than the characteristic (streamwise) wavelength of the primary instability. Such domains allow many of the (essentially local) secondary instabilities to develop and hence trigger turbulence transition. However, as can be straightforwardly observed in sufficiently wide tilting tank experiments (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and sufficiently wide computational domains, as clearly demonstrated by Fritts et al. (2022a, b), inherently three-dimensional ‘knots’, ‘tubes’, and billow ‘defects’ can develop in the spanwise direction on scales of the order of (but typically larger than) the primary instability’s most unstable streamwise wavelength.
Moreover, there will always be an inevitable decay of these reported flow structures unless the underlying driving mechanism (i.e. the shear) is replenished. Although such initial-value-problem mixing events are undoubtedly of geophysical interest and indeed demonstrate the formation of transient larger-scale dynamics exhibiting the so-called ‘lifecycle’ of a shear-driven mixing event, it is important to remember that such initial-value-problem mixing events are really just one end member class of the possible flows of geophysical interest.
@@
The other obvious end member class is the class of continuously forced flows, where the ‘background’ velocity and density (or equivalently buoyancy) profiles are driven by some external, quasi-steady forcing. Possible candidate mechanisms for such geophysically relevant forcings include wind, solar radiation and resulting evaporation at the surface of the ocean (Thorpe, 2005), tidal forcing (Laurent et al., 2002), active matter living in water (Castro et al., 2022), or continuous outflows from rivers (Uncles & Mitchell, 2011). While this list is not meant to be complete in any sense, it illustrates that quasi-steady forcings do occur in geophysically relevant situations. An (artificial) volumetric forcing is a particularly convenient way to mimic these natural complex mechanisms in a simplified manner. As visualised on the right side of figure 1, such a forcing may be defined to relax the time-dependent local profiles.
@@
Such a forcing was introduced by Smith et al. (2021), who demonstrated that – after an initial transient where primary instabilities (either KHI or HWI) develop and (inevitably) break down – the ensuing turbulence could be sustained over arbitrarily long times, thus enabling statistically steady statistics of the flow to be calculated. There is a clear attraction to considering such forced flows due principally to the inherent removal of the confounding effects of the life cycle of the mixing event from the turbulence statistics. Therefore, such flows seem the natural test-bed to study the emergent and sustained turbulent self-organisation of the flow on large scales. This endeavour, however, comes necessarily with the challenge of first diagnosing whether or not there is actually an inherent convergence of the resulting depth of the turbulent shear layer, and second identifying its minimal required extent to enable such statistical convergence (that is unbiased by the extent of the domain).
Given that Smith et al. (2021), similar to the initial value problem flows discussed above, also considered relatively small computational domains only, it is not at all clear whether or not their statistically stationary flow remained unaffected by the computational domain size. An equivalent question to ask is ‘what are the unconstrained emergent length scales of a forced stratified shear flow?’ Answering that (open) question is the central objective of this paper.
@@
To address this objective, the rest of the paper is organised as follows. In section 2 we present our numerical approach before we list the range of considered control parameters and computational domains at the start of section 3. For simplicity, we focus on one set of parameters that enables a relatively weakly stratified flow which is prone to primary KHI. After briefly identifying some interesting early-time dynamics that may possibly be affected by molecular diffusion due to the choice of control parameters (but is not of principal interest here), we shift our focus to the statistically steady turbulent state that is sustained by the volumetric forcing. We identify emergent, strongly anisotropic length scales which are proven to converge in sufficiently extended domains. We demonstrate conclusively that key statistics of the flow, including in particular those related to irreversible mixing, are sensitive to the large-scale self-organisation of the flow or equivalently the extent of the computational domain. From a practical perspective, we remark that for convergent statistics, perhaps surprisingly, the flow domain may need to be extraordinarily extended (i.e. of order $100$ times larger) compared to the initial shear-layer half-depth. We discuss the implications of the discovered emergent large-scale dynamics with an emphasis on the flow physics in section 4 and propose their dynamical origin. Finally, we draw our conclusions in section 5.
@@ 2.1 Governing equations
where $t_{\textrm{r}}$ is the response time while $u_{x,0}$ and $b_{0}$ represent the initial streamwise velocity and buoyancy base profiles to which the flow relaxes back under the effect of the forcing, at least in principle. In this context, $R_{0}:=d_{0}/\delta_{0}$ defines the ratio of initial interface half-depths (i.e. $\delta_{0}$ represents the dimensional initial buoyancy interface half-depth) with $R_{0}=\sqrt{\mathrm{Pr}}$ following the diffusive arguments presented by Smyth et al. (1988). In essence, these forcing terms are idealized models of geophysically relevant processes (see again our introduction) that tend to restore the initial profiles $u_{x,0}$ and $b_{0}$, and thus sustain turbulence over arbitrarily long times. In this context, we stress that the response time $t_{\textrm{r}}$ quantifies the relative strength of these superposed processes. While $t_{\textrm{r}}\rightarrow 0$ implies vigorously enforcing the sharp initial profiles, $t_{\textrm{r}}\rightarrow\infty$ corresponds to the disappearance of the forcing. As shown by Smith et al. (2021), the broad sweetspot – to allow for sustained turbulence without prescribing a significant signal of the forcing – lies in between and tends to preserve the primary instability associated with the underlying base profiles. Finally, we remark that, although the included base profiles are time-independent, the resulting rate of momentum or buoyancy injection is typically not constant in space or time.
Hence, the governing equations (2) – (4) are fully specified via four control parameters: $\mathrm{Pr}$, $\mathrm{Re}_{0}$, $\mathrm{Ri}_{0}$, and $t_{\textrm{r}}$.
@@ 2.2 Domain, boundary and initial conditions, and numerical code
As indicated by figure 1, the volume $V:=L_{x}\times L_{y}\times L_{z}$ of the numerical domain, i.e. the product of the streamwise, spanwise, and vertical extents $L_{x}$, $L_{y}$, and $L_{z}$, respectively. Both the midpoint of the shear layer and the midpoint of the buoyancy distribution are located at midplane, $z=0$, with a horizontal cross-section $A:=L_{x}\times L_{y}$. We consider a horizontally periodic domain where any quantity
@@
The random fluctuations $10^{-3}\leq\Upsilon\left(\bm{x}\right)\leq 10^{-3}$ ‘seed’ the development of primary instabilities.
We solve the coupled governing equations (2) – (4), subject to these boundary and initial conditions (8) – (10), using the GPU-accelerated spectral element solver NekRS (Fischer, 1997; Scheel et al., 2013; Fischer et al., 2022). As the spatial resolution of spectral element methods is determined by the combination of the number of spectral elements $N_{\textrm{e}}=N_{\mathrm{e},x}\times N_{\mathrm{e},y}\times N_{\mathrm{e},z}$ and polynomial order $N$ of the spectral expansion within each element (Deville et al., 2002; Vieweg, 2023), such methods can – as shown in more detail in appendix A – accommodate different required spatial resolutions across the domain and are thus well suited to resolve shear flows efficiently. This is particularly important given degrees of freedom of up to $N_{\textrm{dof}}\approx 3N_{\textrm{e}}N^{3}\sim\mathcal{O}\left(10^{9}\right)$ and the lengthy flow evolution in our present study.
@@
$16^{2}$ $0$ $9^{2}$ $0$ $2.66\pm 0.16$ $2.80\pm 0.14$ $0.95\pm 0.04$ $133\pm\hphantom{0}8$ $0.0333\pm 0.0020$
$32^{2}$ $0$ $18^{2}$ $0$ $35\pm 0.34 5.5$$54\pm 0.35$ $0.96\pm 0.02$ $267\pm 17$ $0.0668\pm 0.0043$
$64^{2}$ $0$ $36^{2}$ $0$ $55\pm 0.47 7.7$$75\pm 0.36$ $0.97\pm 0.03$ $377\pm 24$ $0.0943\pm 0.0059$
$96^{2}$ $0$ $54^{2}$ $0$ $7.92\pm 0.53$ $8.27\pm 0.56$ $0.96\pm 0.02$ $396\pm 26$ $0.0990\pm 0.0066$
$128^{2}$ $0$ $72^{2}$ $0$ $8.05\pm 0.42$ $8.41\pm 0.45$ $0.96\pm 0.01$ $403\pm 21$ $0.1007\pm 0.0053$
$256^{2}$ $0$ $144^{2}$ $0$ $7.90\pm 0.18$ $8.22\pm 0.17$ $0.96\pm 0.01$ $395\pm\hphantom{0}9$ $0.0987\pm 0.0022$
$512^{2}$ $0$ $288^{2}$ $0$ $7.73\pm 0.08$ $8.06\pm 0.07$ $0.96\pm 0.00$ $387\pm\hphantom{0}4$ $0.0967\pm 0.0010$
@@
Table 1: Simulation parameters. The Prandtl number $\mathrm{Pr}=1$, initial bulk Reynolds number $\mathrm{Re}_{0}=50$, initial bulk Richardson number $\mathrm{Ri}_{0}=0.0125$, response time $t_{\textrm{r}}=100$, and initial ratio of interface half-depths $R_{0}=1$ in a horizontally periodic domain. In the vertical direction, the domain has an aspect ratio $L_{z}=48$ spanned by $N_{\mathrm{e},z}=18$ non-uniformly distributed spectral elements (see appendix A for more details) together with free-slip and no-flux boundary conditions for the velocity and buoyancy field, respectively. The polynomial order $N=6$. Although the total evolution or run time of each flow $t_{\textrm{evo}}=5,040$, this work focuses on the statistically stationary dynamics during the last $\Delta t=3,000$ only. For each simulation, this table lists the horizontal cross-sectional areas $L_{x}\times L_{y}$ and corresponding numbers of uniformly distributed spectral elements $N_{\mathrm{e},x}\times N_{\mathrm{e},y}$. Moreover, we include the final emergent shear layer half-depth $d$ of the streamwise velocity field as well as the buoyancy interface half-depth $\delta$, their ratio $R$, the bulk Reynolds number $\mathrm{Re}$, as well as the bulk Richardson number $\mathrm{Ri}$, listing both their temporal means and standard deviations.